OpenStudy (anonymous):

The total area of a cylinder is 100(pie). If r=h, find r.

OpenStudy (anonymous):

The formula to find the total area is L.A + (2)Area of the base.

OpenStudy (anonymous):

total area of cylinder = area of top + area of bottom + area of curved surface \[100\pi=\pi r^2+\pi r^2+2\pi rh\\ 100\pi=4\pi r^2\]

OpenStudy (anonymous):

So the answer is: \[4xr ^{2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

solve for "r"

OpenStudy (anonymous):

the question says, "find r"